### v1

# Training and Evaluating with Human Label Variation: An Empirical Study

Abstract
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth, instead embracing the natural variation in human labelling to train and evaluate models. While various training methods and metrics for HLV have been proposed, there has been no systematic meta-evaluation of HLV evaluation metrics, contributing to the lack of clarity in the best HLV training method. We propose new evaluation metrics and training methods and empirically meta-evaluate HLV evaluation metrics. We find that training on either disaggregated annotations or soft labels often performs best across metrics, and that our proposed soft metric correlates best with human preference. [^1] Code is available at: https://github.com/kmkurn/train-eval-hlv
1 Introduction
Human label variation (HLV) challenges the standard assumption that an example has a single ground truth (Plank, 2022). Without this assumption, model training and evaluation are no longer straightforward. To address this issue, recent work has proposed various methods to train models and metrics to evaluate their performance in the HLV context. However, there has not been a systematic meta-evaluation of HLV evaluation metrics, resulting in prior work reporting model performance across a wide variety of metrics, thwarting systematic model comparison. This has in turn contributed to a lack of clarity in the best HLV training method (Uma et al., 2021). In addition, NLP has largely ignored the field of remote sensing that has dealt with issues in model evaluation against a fuzzy ground truth.
In this work, we: (a) propose soft evaluation metrics, taking inspiration from remote sensing research; (b) propose new training methods based on the soft evaluation metrics; and (c) perform an empirical meta-evaluation of HLV evaluation metrics. Most existing metrics for HLV represent human judgements as probability distributions over classes and employ information-theoretic measures such as divergence. Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. To the best of our knowledge, we are the first to propose such soft metrics for HLV.
Because they are differentiable, we then use these soft metrics as training objectives. Experiments on 6 datasets show that training on either disaggregated annotations or soft labels often achieves the best performance across metrics. To perform the meta-evaluation, we recruit expert annotators and collect their preference on predicted judgement distributions. The meta-evaluation shows that our proposed soft metric correlates best with human preference, but the correlation is modest, highlighting the need for better evaluation metrics for HLV.
2 Related Work
Training methods for human label variation
First introduced by Plank (2022), the term human label variation (HLV) captures the fact that disagreeements in annotations can be well-founded and thus signal for data-driven methods. It challenges the traditional notion in machine learning that an instance has a single ground truth. Training methods that can handle such variation had been proposed before the term was coined (Sheng et al., 2008; Peterson et al., 2019; Uma et al., 2020, inter alia). Newer methods have also been employed in the Learning with Disagreement shared task (Leonardelli et al., 2023) which provides a benchmark for HLV. While a systematic investigation of these methods exists (Uma et al., 2020), it employs smaller pretrained language models and covers only binary or multiclass classification tasks. In contrast, our work additionally employs a large language model and covers multilabel tasks. More importantly, our work is the first to empirically meta-evaluate commonly-used evaluation metrics in the HLV context. In addition, we propose new soft metrics, which we also include in the meta-evaluation.
Evaluation of soft classification in remote sensing
While evaluating in the HLV context is a relatively new concept in NLP, evaluating against a fuzzy [^2] The terms “fuzzy” and “crisp” are sometimes used in remote sensing literature to mean “soft” and “hard” in NLP literature. reference is a well-studied area in remote sensing research. Early work by Foody (1996) used a measure of distance that is equivalent to twice the Jenson-Shannon divergence (Lin, 1991). Binaghi et al. (1999) argued that entropy-based measures are sensible only if the reference is crisp, and proposed the use of fuzzy set theory to compute the soft version of standard evaluation metrics such as accuracy. A similar approach was also proposed by Harju and Mesaros (2023) for audio data. Key to this approach is the minimum function used to compute the class membership scores given a fuzzy reference and a fuzzy model output. Other proposed functions include product (Lewis and Brown, 2001), sum (Pontius and Connors, 2006), and a composite of such functions designed to ensure diagonality of the confusion matrix when the reference and the model output match perfectly (Pontius and Cheuk, 2006; Silván-Cárdenas and Wang, 2008). Approaches that take the cost of misclassifications into account have also been proposed (Gómez et al., 2008). In this work, we take the fuzzy set approach by Binaghi et al. (1999) and apply it to text classification tasks.
3 Evaluation Metrics for HLV
Table 1 provides an overview of evaluation metrics used in this work.
Notation
Let $P_{ik}\geq 0$ denote the true human judgement for example $i$ and class $k$, i.e. proportion of humans labelling example $i$ with class $k$. Let $Q_{ik}\geq 0$ denote the value of $P_{ik}$ predicted by a model. An evaluation metric $m$ is a function such that the scalar $m(\mathbf{P},\mathbf{Q})$ measures how well (i.e., a positive orientation) $\mathbf{Q}$ captures the judgement distribution in $\mathbf{P}$. To complete our notation, let $N$ and $K$ denote the number of examples and classes respectively.
Multiclass Multilabel
Existing: Accuracy Micro F1
Macro F1 Macro F1
PO-JSD
Entropy correlation
Proposed: Soft accuracy Soft micro F1
Soft macro F1 Soft macro F1
Multilabel PO-JSD
Multilabel entropy correlation
Table 1: HLV metrics for classification tasks.
3.1 Multiclass Metrics
In multiclass classification tasks, an example can only be assigned to one class. This is reflected by the unity constraints $\sum_{k}P_{ik}=1$ and $\sum_{k}Q_{ik}=1$ for all $i$. We study a total of 6 multiclass evaluation metrics $m$. The first two are standard hard metrics, the next two are their soft versions that we propose inspired by prior work (Binaghi et al., 1999; Harju and Mesaros, 2023), and the last two are existing HLV metrics.
Let $I^{P}_{ik}$ and $I^{Q}_{ik}$ denote $\mathds{1}(\arg\max_{l}P_{il}=k)$ and $\mathds{1}(\arg\max_{l}Q_{il}=k)$ respectively. The 6 metric definitions under our notation are as follows:
  1. 1.
    (Hard) accuracy, where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{N}\sum_{ik}I^{P}_{ik}I^{Q}_{ik}.$
    This is the standard evaluation metric for classification tasks.
  2. 2.
    (Hard) macro F1, where
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{K}\sum_{k}\frac{2\sum_{i}I^{P}_{ik}I^{Q}_{ik}}{\sum_{i}\left(I^{P}_{ik}+I^{Q}_{ik}\right)}.$
    This addresses the issue with accuracy that it is biased toward the majority class when class proportions are imbalanced.
  3. 3.
    Soft accuracy, where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{N}\sum_{ik}\min(P_{ik},Q_{ik}).$
    Novel to this work, for evaluation in the context of HLV. This is a special case of the soft micro F1 score proposed later in Section 3.2.
  4. 4.
    Soft macro F1, where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{K}\sum_{k}\frac{2\sum_{i}\min(P_{ik},Q_{ik})}{\sum_{i}(P_{ik}+Q_{ik})}.$
    Also novel to this work, to address the issue of class imbalance with soft accuracy and other existing HLV metrics. [^3] In the context of HLV, we understand the concept of class imbalance analogously to that in the context where class assignments are hard: there exists a class $k$ such that $\sum_{i}P_{ik}\operatorname{>\!\!>}\sum_{i}P_{il}$ for all $l\neq k$. Note that this metric implies the existence of soft class-wise metrics (see Appendix A).
  5. 5.
    Jensen-Shannon divergence (Uma et al., 2021), which we modify into its positvely oriented version (PO-JSD for short) where:
    $m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}1-\frac{1}{N}\sum_{i}\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i}).$
    Vectors $\mathbf{p}_{i},\mathbf{q}_{i}$ denote the $i$-th row of $\mathbf{P},\mathbf{Q}$ respectively. The scalar $\operatorname{JSD}(\mathbf{a},\mathbf{b})=\frac{1}{2}\left(\operatorname{KL}(\mathbf{a},\frac{1}{2}\left(\mathbf{a}+\mathbf{b}\right))+\operatorname{KL}(\mathbf{b},\frac{1}{2}\left(\mathbf{a}+\mathbf{b}\right))\right)$ is the Jensen-Shannon divergence (Lin, 1991), where $\operatorname{KL}(\mathbf{a},\mathbf{b})=\sum_{k}a_{k}\log_{2}\frac{a_{k}}{b_{k}}$ is the Kullback-Leibler divergence. [^4] Jensen-Shannon divergence of two distributions has an upper bound of $\log_{b}2$ if the logarithms used in $\operatorname{KL}(\cdot)$ are of base $b$. Normalising this bound to 1 results in logarithms of base 2 as $\log_{b}x/\log_{b}2=\log_{2}x$.
  6. 6.
    Entropy correlation (Uma et al., 2020), where:
    $m(\mathbf{P},\mathbf{Q})$ =$\overset{\mathrm{def}}{=}\frac{\sum_{i}\zeta^{P}_{i}\zeta^{Q}_{i}}{\sqrt{\sum_{i}{\zeta^{P}_{i}}^{2}}\sqrt{\sum_{i}{\zeta^{Q}_{i}}^{2}}},$
    $\zeta^{P}_{i}$ $=\eta^{P}_{i}-\bar{\eta}^{P},\text{ and }$
    $\zeta^{Q}_{i}$ $=\eta^{Q}_{i}-\bar{\eta}^{Q}.$
    Scalar $\eta^{P}_{i}=-\sum_{k}\frac{P_{ik}\log P_{ik}}{\log K}$ is the normalised entropy of row $i$ of $\mathbf{P}$, and $\bar{\eta}^{P}=\frac{1}{N}\sum_{i}\eta^{P}_{i}$ is the mean of $\{\eta^{P}_{i}\}_{i=1}^{N}$. Scalars $\eta^{Q}_{i}$ and $\bar{\eta}^{Q}$ are defined analogously.
By noting that $N=\sum_{ik}P_{ik}$, soft accuracy can be interpreted as the proportion of the judgement distribution that is correctly predicted. This interpretation is similar to the hard counterpart (i.e. proportion of correctly predicted examples) and arguably more intuitive than that of existing metrics because it is stated directly in terms of the problem of predicting human judgement distribution. Other similarities include symmetry (due to the symmetry of $\min$) and boundedness (between 0 and 1), where the lower and the upper bounds are achieved if and only if $P_{ik}Q_{ik}=0$ and $P_{ik}=Q_{ik}$ respectively for all $i,k$. Note also that soft accuracy is reduced to the standard hard accuracy when all rows of both $\mathbf{P}$ and $\mathbf{Q}$ are one-hot vectors (i.e. no label variation).
While functions other than $\min$ have been used to compute soft accuracy (see Section 2), they lead to the soft accuracy being $<1$ in the case where $\mathbf{P}=\mathbf{Q}$. This is problematic because intuitively, good HLV evaluation metrics should produce the highest score when the predicted judgement distribution perfectly matches the true counterpart. Therefore, we use the $\min$ function in this work.
Other possible evaluation metrics include entropy similarity (Uma et al., 2021), (negative) cross-entropy (Peterson et al., 2019; Pavlick and Kwiatkowski, 2019), (negative) Kullback-Leibler divergence (Nie et al., 2020), (the positively oriented version of) information closeness (Foody, 1996) and (the positively oriented version of) Jensen-Shannon distance (Nie et al., 2020). However, they are excluded from this study because of their expected high correlation with either PO-JSD or entropy correlation.
3.2 Multilabel Metrics
Next, we consider the case where the unity constraints do not hold, i.e. an example can be (soft) assigned to multiple classes, forming a multilabel classification task.
By considering this case as multiple independent binary classification tasks, some existing metrics can be extended straightforwardly by taking the average of the metric values over classes. For example, PO-JSD can be modified into the multilabel version:
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}1-\frac{1}{NK}\sum_{ik}\operatorname{JSD}(P^{\text{bin}}_{ik},Q^{\text{bin}}_{ik})$
where $P^{\text{bin}}_{ik}=[P_{ik},1-P_{ik}]$ is a judgement distribution over two classes, and $Q^{\text{bin}}_{ik}$ is defined analogously. Entropy correlation can also be modified similarly into the multilabel version:
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}\frac{1}{K}\sum_{k}\frac{\sum_{i}\zeta^{P}_{ik}\zeta^{Q}_{ik}}{\sqrt{\sum_{i}{\zeta^{P}_{ik}}^{2}}\sqrt{\sum_{i}{\zeta^{Q}_{ik}}^{2}}}$
where $\zeta^{P}_{ik}=\eta^{P}_{ik}-\bar{\eta}^{P}_{k}$, $\eta^{P}_{ik}=-\frac{1}{\log 2}\left(P_{ik}\log P_{ik}+(1-P_{ik})\log(1-P_{ik})\right)$, $\bar{\eta}^{P}_{k}=\frac{1}{N}\sum_{i}\eta^{P}_{ik}$, and scalars $\zeta^{Q}_{ik}$, $\eta^{Q}_{ik}$, and $\bar{\eta}^{Q}_{k}$ are defined analogously.
While it is possible to define (hard) accuracy for this multilabel case, a more common evaluation metric is the micro F1 score, which in our notation can be expressed as:
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}2\frac{\sum_{ik}J^{P}_{ik}J^{Q}_{ik}}{\sum_{ik}\left(J^{P}_{ik}+J^{Q}_{ik}\right)}$
where $J^{P}_{ik}=\mathds{1}(P_{ik}>0.5)$, and $J^{Q}_{ik}$ is defined analogously. Similarly, we define the soft version of the micro F1 score as follows (see derivations in Appendix A):
$m(\mathbf{P},\mathbf{Q})\overset{\mathrm{def}}{=}2\frac{\sum_{ik}\min(P_{ik},Q_{ik})}{\sum_{ik}(P_{ik}+Q_{ik})}.$
The (hard) micro F1 score is a special case of this soft counterpart when $P_{ik}\in\{0,1\}$ and $Q_{ik}\in\{0,1\}$ (i.e. no label variation). Furthermore, the soft accuracy in Section 3.1 is a special case of this soft micro F1 score when the unity constraints hold. Therefore, the soft micro F1 score is a general evaluation metric for classification tasks.
$K$ A-J A-E A-S J-E J-S E-S
$\alpha=\beta=10$
10 0.272 3470.2723470.272\,3470.272 347 0.050 2550.0502550.050\,2550.050 255 0.289 9840.2899840.289\,9840.289 984 0.035 9270.0359270.035\,9270.035 927 0.942 2630.9422630.942\,2630.942 263 0.100 3030.1003030.100\,3030.100 303
100 0.047 5880.0475880.047\,5880.047 588 0.040 9640.0409640.040\,9640.040 964 0.051 9560.0519560.051\,9560.051 956 0.091 1220.0911220.091\,1220.091 122 0.924 1290.9241290.924\,1290.924 129 0.087 5600.0875600.087\,5600.087 560
$\alpha=\beta=0.1$
10 0.627 7730.6277730.627\,7730.627 773 0.063 8200.0638200.063\,8200.063 820 0.733 7430.7337430.733\,7430.733 743 0.126 1640.1261640.126\,1640.126 164 0.963 2140.9632140.963\,2140.963 214 0.162 8160.1628160.162\,8160.162 816
100 0.194 0690.1940690.194\,0690.194 069 0.093 1130.0931130.093\,1130.093 113 0.251 3550.2513550.251\,3550.251 355 0.014 8690.0148690.014\,8690.014 869 0.961 0860.9610860.961\,0860.961 086 0.032 2250.0322250.032\,2250.032 225
$\alpha=10,\beta=0.1$
10 0.226 1750.2261750.226\,1750.226 175 0.015 882-0.015882-0.015\,882- 0.015 882 0.206 4640.2064640.206\,4640.206 464 0.018 527-0.018527-0.018\,527- 0.018 527 0.975 1440.9751440.975\,1440.975 144 0.002 286-0.002286-0.002\,286- 0.002 286
100 0.004 257-0.004257-0.004\,257- 0.004 257 0.062 204-0.062204-0.062\,204- 0.062 204 0.026 900-0.026900-0.026\,900- 0.026 900 0.050 0800.0500800.050\,0800.050 080 0.957 8160.9578160.957\,8160.957 816 0.061 6400.0616400.061\,6400.061 640
$\alpha=0.1,\beta=10$
10 0.211 8740.2118740.211\,8740.211 874 0.016 4670.0164670.016\,4670.016 467 0.181 7060.1817060.181\,7060.181 706 0.016 632-0.016632-0.016\,632- 0.016 632 0.975 8300.9758300.975\,8300.975 830 0.019 553-0.019553-0.019\,553- 0.019 553
100 0.103 4780.1034780.103\,4780.103 478 0.021 530-0.021530-0.021\,530- 0.021 530 0.057 2680.0572680.057\,2680.057 268 0.005 990-0.005990-0.005\,990- 0.005 990 0.967 4200.9674200.967\,4200.967 420 0.005 4470.0054470.005\,4470.005 447
Table 2: Pearson correlations between a pair of evaluation metrics $m(\mathbf{P},\mathbf{Q})$ for 1K examples and various number of classes ($K$) where the rows of $\mathbf{P}$ and $\mathbf{Q}$ are drawn from a symmetric Dirichlet with parameters $\alpha$ and $\beta$ respectively. A, J, E, and S denote the accuracy, PO-JSD, entropy correlation, and soft accuracy respectively.
3.3 Relationship between Metrics
To analyse the relationship between metrics, we compute Pearson correlation between each pair of multiclass metrics (excluding the macro ones) for various number of classes (see Appendix B for details), similar to prior work (Chicco and Jurman, 2020).
Table 2 shows that the metrics are generally weakly correlated with each other, with the exception of soft accuracy and PO-JSD where the two are consistently strongly correlated ($r>0.9$). This strong correlation with an established HLV metric gives an empirical assurance that our proposed soft accuracy is a sensible metric. The table also shows that soft accuracy is only moderately correlated with (hard) accuracy in most cases, where this correlation gets weaker when the number of classes is large ($K=100$), or either $\mathbf{P}$ or $\mathbf{Q}$ is dense ($\alpha=10$ or $\beta=10$). This finding suggests that hard and soft accuracy capture different aspects when humans disagree. Furthermore, the table shows that entropy correlation is poorly correlated with all metrics, which suggests that it is an outlier HLV metric.
Soft Accuracy and PO-JSD
Figure 1: Soft accuracy and PO-JSD graphs for a binary classification problem on a single example where the true and the predicted judgement for the positive class is 0.5 and $q$ respectively.
Despite their strong positive correlation, soft accuracy and PO-JSD have a notable difference: soft accuracy is upper-bounded by PO-JSD (see Appendix C for proof). This fact implies that soft accuracy penalises a model more heavily than PO-JSD, as illustrated in Figure 1. It shows that a skewed predicted judgement distribution of $(0.2,0.8)$ has a soft accuracy of $0.7$, much less than the PO-JSD ($>0.9$). The high PO-JSD is potentially misleading as it is very close to the maximum possible value of 1.
4 Training Methods for HLV
We experiment with a total of 14 HLV training methods, as summarised in Table 3. The first method (ReL) is the simplest and uses standard cross-entropy loss as the training objective based on prior work (Sheng et al., 2008). The next 8 methods (MV, AR, ARh, AL, SL, SLMV, AE, AEh) are from the Learning with Disagreements shared task (Leonardelli et al., 2023) which also use standard cross-entropy loss. The next 3 methods (JSD, SmF1, SMF1) use soft metrics that are differentiable as the training objective. The last 2 methods (LA-min, LA-max) aggregate the cross-entropy losses of all annotations for a given input.
Name A M L
Repeated labelling (ReL)
Majority voting (MV)
Annotator ranking (AR)
Annotator ranking (hard) (ARh)
Ambiguous labelling (AL)
Soft labelling (SL)
Multitask of SL and MV (SLMV)
Annotator ensemble (AE)
Annotator ensemble (hard) (AEh)
Below are proposed in this work
Jensen-Shannon divergence (JSD)
Soft micro F1 (SmF1)
Soft macro F1 (SMF1)
Loss aggregation with $\min$ (LA-min)
Loss aggregation with $\max$ (LA-max)
Table 3: Overview of HLV training methods used, along with details of whether annotator identity is required (A), multilabel tasks are supported (M), and training data is enlarged (L) which makes training run much longer.
Name Labels Language Annotators $N$ $J$ $K$ A M
HS-Brexit (Akhtar et al., 2021) Hate speech English Experts 1K $6.0$ 2
MD-Agreement (Leonardelli et al., 2021) Offensiveness English Crowds 10K $5.0$ 2
ArMIS (Almanea and Poesio, 2022) Misogyny Arabic Experts 1K $3.0$ 2
ChaosNLI (Nie et al., 2020) NLI English Crowds 3K $100.0$ 3
MFRC (Trager et al., 2022) Moral sentiment English Experts 18K $3.4$ 8
Table 4: Public datasets used in the empirical meta-evaluation of metrics, along with the number of examples to the nearest thousand ($N$), average number of annotations per example ($J$), number of classes ($K$), whether annotator identity is present (A), and whether the task is multilabel classification (M).
  1. 1.
    Training on the disaggregated labels directly (Sheng et al., 2008), which we refer to as repeated labelling (ReL). For example, if the training data only has a single instance $x$ with annotations $y_{1}$ and $y_{2}$ then ReL constructs new training data $\{(x,y_{1}),(x,y_{2})\}$.
  2. 2.
    Training on the majority-voted labels (MV). This is the baseline model given by the shared task organisers.
  3. 3.
    Instance weighting, where the weights are the sum of annotator ranking scores for that instance (Cui, 2023). Suppose that training instance $x_{1}$ is labelled as A, A, B, and A by 4 annotators; and $x_{2}$ is labelled as A, B, and B by only the first 3 annotators. There are two ways of computing the ranking score of an annotator:
    1. (a)
      Average judgement of the majority-voted label given by the annotator (AR). The annotator ranking scores are (in order) $\frac{3}{4}$, $\frac{1}{2}\left(\frac{3}{4}+\frac{2}{3}\right)$, $\frac{2}{3}$, and $\frac{3}{4}$.
    2. (b)
      Average number of times the annotator agrees with the majority (ARh). The scores are $\frac{1}{2}$, $\frac{2}{2}$, $\frac{1}{2}$, and $\frac{1}{1}$.
    For both methods, the ranking score is zero if the annotator never agrees with the majority. Majority-voted labels are used as ground truth for training. The training objective is the cross-entropy loss multiplied by the instance weights.
  4. 4.
    Ambiguous labelling (AL) for instances with disagreements (Gajewska, 2023). In short, if annotators are not unanimous in their judgement then the instance is labelled with all selected classes exactly once and included in the training set. Because it was developed for single-label tasks, we exclude this method from multilabel tasks. [^5] Extending this method to multilabel tasks is beyond the scope of this work.
  5. 5.
    Training with soft labels (SL) as targets (Maity et al., 2023; Wan and Badillo-Urquiola, 2023). For example, if an instance is labelled as positive by 80% of annotators then the log-likelihood is $0.8\log p(1)+0.2\log p(0)$, where $p(1)$ and $p(0)$ denote the predicted probability of the positive and the negative class respectively.
  6. 6.
    Multi-task learning using both the majority-voted and the soft labels as targets (SLMV; Grötzinger et al., 2023). Two models with a shared encoder are trained to predict the two types of labels respectively, which is similar to the method of Fornaciari et al. (2021). For example, if an instance is labelled as positive by 80% of annotators then the multitask log-likelihood is $0.8\log p_{\theta_{1}}(1)+0.2\log p_{\theta_{1}}(0)+\log p_{\theta_{2}}(1)$, where $\theta_{1}$ and $\theta_{2}$ denote the parameters of the first and the second model respectively. We predict only the soft labels for evaluation.
  7. 7.
    Modelling each annotator separately followed by ensembling (Sullivan et al., 2023; Vitsakis et al., 2023). In this approach, we have as many models as there are annotators. Each model is trained on the labels given by a distinct annotator. In practice, we share the encoder parameters across models. Suppose that there are 3 annotators, and the predicted judgements of the positive class for an instance are 60%, 30%, and 90%. At test time, there are 2 ways to ensemble these predictions:
    1. (a)
      Taking the average (AE; Sullivan et al., 2023). The final predicted judgement for the positive class is $\frac{1}{3}\left(0.6+0.3+0.9\right)$.
    2. (b)
      Computing the distribution of most likely labels (AEh; Vitsakis et al., 2023). The final predicted judgement for the positive class is $\frac{2}{3}$ because 2 out of 3 judgements are over 50%.
  8. 8.
    Training using JSD (JSD), $1-$ soft accuracy for multiclass or micro F1 for multilabel tasks (SmF1), and $1-$ soft macro F1 (SMF1) as the objective (see Section 3 for definitions). [^6] We exclude entropy correlation because it doesn’t have a unique maximiser. These approaches leverage the differentiability of soft metrics, resulting in a single objective for both training and inference. By directly optimising the evaluation metric at training time, we expect the models to exhibit superior performance, especially when evaluated with the corresponding metric.
  9. 9.
    Aggregating the losses of all annotations for a given training instance. For example, given a training instance $x$ with annotations $y_{1}$ and $y_{2}$, we compute the losses $l_{1}=\mathcal{L}(x,y_{1})$ and $l_{2}=\mathcal{L}(x,y_{2})$ where $\mathcal{L}$ is the cross-entropy loss function. Then, we update model parameters based on the gradient of $g(l_{1},l_{2})$ where $g$ is an aggregation function. We experiment with the $\min$ (LA-min) and the $\max$ (LA-max) functions as $g$. Using the $\min$ (resp. $\max$) function means selecting the least (resp. most) “surprising” annotation for the model. Using the mean aggregation is mathematically equivalent to SL and thus excluded.
5 Evaluation of HLV Training Methods
We now investigate if there is an HLV training method that consistently performs well across all datasets and evaluation metrics.
5.1 Experimental Setup
Public datasets
We use 5 public datasets with disaggregated annotations that have been used in previous HLV studies, four of which are in English and one in Arabic. Table 4 shows that the datasets cover both crowd (untrained) and expert (trained) annotators, as well as numbers of examples and annotations, and the language of the dataset. For HS-Brexit, MD-Agreement, and ArMIS, we use the same train–test splits as Leonardelli et al. (2023). For both ChaosNLI and MFRC, we use 10-fold cross-validation with 10% of the training portion used as a development set. For ChaosNLI, we include only the SNLI and MNLI subsets (each has 1.5K examples) because of their standard NLI setup of one premise and one hypothesis.
Text Annotation Game dataset
We also include a dataset called TAG which consists of English texts describing legal problems written by non-expert legal help seekers. We have access to this dataset because of our collaboration with Justice Connect, [^7] https://justiceconnect.org.au an Australian public benevolent institution [^8] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile providing free legal assistance to laypeople facing legal problems. Each text in the dataset was annotated by an average of 5.5 practising lawyers, who each selected one or more out of the 33 areas of law [^9] Including a special label “Not a legal issue”. that applied to the problem. Thus, it is a multilabel classification task. Example areas of law include Neighbourhood disputes and Housing and residential tenancies. Average inter-annotator agreement ($\alpha$) over the areas of law is $0.454$, which is modest. [^10] We compute this agreement on random 10% selection of the data due to the high computational cost. This is because lawyers often have different interpretations of the same problem due to their different legal specialisations and years of experience. However, these different interpretations are all valid as human label variation because the lawyers are subject matter experts. The dataset has a total of 11K texts collected between July 2020 and early December 2023. We randomly split the dataset 8:1:1 for the training, development, and test sets respectively. While it cannot be released publicly because it is based on real-world confidential legal requests from help-seekers, we include the TAG dataset due to its importance for the meta-evaluation later in Section 6.
Models
We experiment with both small and large language models. Because HS-Brexit, MD-Agreement, and ArMIS use Twitter as source data, we use TwHIN-BERT (Zhang et al., 2023) which was pretrained on Twitter data and supports both English and Arabic. For other datasets, we use RoBERTa (Liu et al., 2019). In addition, we experiment with 8B LLaMA 3 (Grattafiori et al., 2024) for the English datasets. [^11] While LLaMA 3 was trained on multilingual data, its intended use is English only. We replace the output layer of these pretrained models with a linear layer with $K$ output units. We use the base versions of both RoBERTa and TwHIN-BERT (with roughly 100M parameters each).
Training
[…]
Evaluation
When cross-validation is used, we evaluate the concatenated test predictions. Otherwise, we evaluate on the test set. We perform the evaluation across 3 training runs.
Figure 2: Performance difference (delta) with mean MV performance on HS-Brexit, MD-Agreement, the SNLI portion of ChaosNLI (C-SNLI), MFRC, and TAG datasets. Both ArMIS and the MNLI portion of ChaosNLI are omitted because the trend is similar (their results are reported in Figure 4 in the appendix). BERT denotes either TwHIN-BERT or RoBERTa, depending on the dataset. C-SNLI doesn’t have annotator identity information, so AE, AEh, AR, ARh are inapplicable. AL is omitted from MFRC and TAG because of its non-trivial application to the multilabel case. SmF1 with TwHIN-BERT predicts zero for all test instances on HS-Brexit, while AEh with RoBERTa does so on several classes in TAG, so their entropy correlations are undefined. The methods are sorted by their mean ranking across datasets, models, and metrics.
5.2 Results
We compute the difference between the performance of a single run of an HLV training method and the mean performance of MV. We then report the mean differences across runs in Figure 2 (performance of MV is reported in Table 7 in the appendix). We make the following observations.
First, HLV training methods have different effectiveness compared to MV, with some methods performing very poorly. For example, on the SNLI portion of ChaosNLI, LA-min consistently performs worse than MV. This finding suggests the importance of choosing the right training method when embracing HLV.
Second, ReL and SL substantially outperform MV on most datasets across models and evaluation metrics. Furthermore, ReL and SL are often among the top-performing methods, even when compared to JSD and SmF1 evaluated using PO-JSD and soft accuracy/micro F1 respectively. In other words, ReL and SL are often on par with (or outperform in some cases with LLaMA) the training method whose learning objective is the same as the evaluation metric, demonstrating the strength of ReL and SL.
Third, ReL and SL are competitive with each other. However, on both HS-Brexit and the SNLI portion of ChaosNLI, ReL outperforms SL with LLaMA. This finding suggests that ReL is better suited to large language models than SL.
Fourth, SMF1 and JSD generally also perform well, although they lag behind ReL and SL. Because both SMF1 and JSD use a soft evaluation metric as the training objective, this finding underlines the value in using differentiable metrics for HLV model training.
We also observe that SmF1 with TwHIN-BERT results in undefined entropy correlation on HS-Brexit. Looking closely, we find that the method degenerates to predicting $Q_{ik}=0$ for all $i$. We observe the same degenerate results for AEh with RoBERTa on the TAG dataset. Moreover, SmF1 with RoBERTa attains much lower soft micro F1 score than the top-performing ReL on TAG, even though SmF1 optimises the evaluation metric directly during training. Our investigation suggests that these unexpected observations are due to the class imbalance present in both HS-Brexit and TAG. When the training objective is fair (e.g., SMF1), the problem disappears.
5.3 Discussion
That ReL and SL perform best is interesting because they are perhaps two of the most straightforward methods to incorporate HLV in model training. This finding underscores that even baseline HLV training methods have high utility. Furthermore, our finding that SL often performs best is in agreement with that of Uma et al. (2021).
That ReL performs well shows that models are capable of learning from disagreeing annotations directly, in contrast with prior work (Uma et al., 2021; Kurniawan et al., 2024). Moreover, while the results suggest that this capability isn’t limited to large language models, ReL is likely to outperform SL when such models are used. We hypothesise that this is due to the enhanced learning capacity of such models, allowing them to learn from disagreeing annotations more flexibly. We leave the investigation of this effect to future work.
Despite its strengths, ReL has substantial computational cost: because it keeps the annotations disaggregated, the size of the training data can be enormous. For example, in the ChaosNLI dataset, ReL makes the training data 100 times larger because there are exactly 100 annotations for each training instance.
6 Meta-Evaluation of HLV Evaluation Metrics
While we have determined effective HLV training methods based on performance across metrics, we haven’t yet identified which metric is the best. Intuitively, a good evaluation metric would correlate well with human preference. We now perform a meta-evaluation of the HLV evaluation metrics based on this intuition.
To obtain a reliable estimate of human preference, access to subject matter experts is essential. Therefore, when they disagree, we can be confident that the disagreement is well-founded rather than due to a misunderstanding or similar noise. In our case, we have access to qualified practising lawyers annotating the TAG dataset whose qualification is standardised and directly related to the annotation task (i.e. legal area classification). Therefore, we use the TAG dataset for this meta-evaluation despite the inability to distribute it.
6.1 Experimental Setup
Figure 3: Relationships between evaluation metrics and the fraction of times humans select the predicted judgement distribution on the TAG dataset. The shaded area denotes a 95% confidence interval. Each data point corresponds to a specific pairing of a language model (e.g., LLaMA) and a training method (e.g., ReL). AEh with RoBERTa is excluded because of degenerate results on entropy correlation.
For each pair of pretrained model and training method, we randomly sample an average of 28 test instances from TAG. For each instance, we (a) show the instance text and both the true and the predicted judgement distributions and (b) ask lawyers which judgement distribution [^14] Because the task is multilabel, the judgement distribution doesn’t necessarily sum to one. is more accurate. The lawyers can choose exactly one, both, or neither of them. To encourage the lawyers to form their own expectation on the judgment distribution, we first show only the instance text and ask them to annotate it with areas of law (see Figure 5 in the appendix for the annotation interface). The mean number of lawyers is 3.8, with an overall inter-annotator agreement ($\alpha$) of $0.167$. The low agreement is expected and not due to noise because the task is inherently subjective, and these annotators are practising lawyers who draw on their real-world legal experience rather than tie-breaking strategies in third-party annotation guidelines.
6.2 Results
We meta-evaluate HLV evaluation metrics by computing the correlation between their value and the fraction of times the lawyers select the predicted judgement distribution. Figure 3 shows that all metrics exhibit a linear relationship with human preference, which is positive for all metrics other than entropy correlation, suggesting that it is a poor HLV metric.
Table 5 confirms that all metrics are positively correlated with human preference except for entropy correlation. Furthermore, it shows that only soft micro F1 and PO-JSD correlate with human judgements at a level of statistical significance. This finding suggests that both metrics are good for evaluation in the HLV context. However, none of the metrics have significant Spearman correlation coefficient.
[…]
Type Metric Pearson $r$ Spearman $\rho$
Hard Micro F1 $0.442$ $0.335$
Macro F1 $0.318$ $0.271$
Soft Entropy correlation $-0.196$ $-0.338$
PO-JSD $0.692$ $0.272$
Soft micro F1 $0.611$ $0.379$
Soft macro F1 $0.291$ $0.164$
[…]
6.3 Method ranking based on human preference
Given the human annotations above, we can compute the performance of an HLV training method based on the number of times humans select the predicted judgement distribution. Table 6 shows that the rankings of the methods have some similarities to TAG results in Figure 2. For instance, both ReL and SL are ranked relatively highly, except for SL with LLaMA where it is ranked 9th. JSD is also consistently ranked highly for both RoBERTa and LLaMA, where for the former it is ranked 1st.
However, the table also shows major discrepancies. AR is consistently ranked highly, even 1st for LLaMA, but this is not at all reflected in Figure 2. LA-min is ranked 2nd and 3rd for RoBERTa and LLaMA respectively, but is often the worst method for a given dataset. That said, these discrepancies are expected because the HLV evaluation metrics only have moderate correlation with human preference (see Table 5).
6.4 Discussion
Our findings suggest that both PO-JSD and soft micro F1 are reasonably good evaluation metrics in the HLV context. Therefore, this finding provides an empirical answer to the question of the best way to perform evaluation when disagreements abound, as raised by Basile et al. (2021).
RoBERTa LLaMA
Rank 1–7 Rank 8–13 Rank 1–7 Rank 8–13
JSD SLMV AR MV
LA-min SMF1 SmF1 SL
SL ARh LA-min ARh
MV AE ReL AEh
AR LA-max JSD AE
ReL AEh SLMV SMF1
SmF1 LA-max
Table 6: Rankings of HLV training methods based on human preference on TAG. With RoBERTa, AEh is ranked last due to its degenerate results.
Uma et al. (2021) pointed out that the value of PO-JSD is confined within a small range. Furthermore, PO-JSD lacks an interpretation accessible to lay people. In contrast, soft micro F1 has both an intuitive interpretation and higher Spearman correlation with human preference. Therefore, we recommend to report both metrics, but select soft micro F1 score when an accessible interpretation is important (e.g., communicating with non-technical people) or one is restricted to a single metric (e.g., in hyperparameter tuning).
7 Conclusions
We propose new training methods and evaluation metrics for human label variation (HLV). Experiments across 14 methods, 6 metrics, and 7 datasets covering both multilabel tasks and expert annotations show that training on disaggregated annotations (ReL) or soft labels (SL) performs best across evaluation metrics. Leveraging the fuzzy set theory, our proposed soft metrics have an intuitive interpretation and are reduced to standard hard metrics if there is no label variation. Our analysis shows that our proposed soft metric is weakly correlated with most existing metrics, providing unique insights into model performance. Our meta-evaluation of HLV evaluation metrics show that the proposed soft metric correlates best with human preference. However, the correlation is modest, highlighting the need for better HLV evaluation metrics.
Limitations
The TAG dataset is private and cannot be released publicly, which is a major limitation of our work in terms of reproducibility. This is because the data is based on real-world confidential legal requests from help-seekers. Nevertheless, we believe our work still offers valuable scientific knowledge on the investigation of HLV training methods and evaluation metrics.
Our empirical meta-evaluation is limited to a multilabel task. Thus, it is unclear if the findings extend to binary or single-label tasks. We leave this empirical meta-evaluation to future work.
Acknowledgements
This research is supported by the Australian Research Council Linkage Project LP210200917151515https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution. [^16] As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface.
Appendix A Derivation of Soft F1 Scores
A.1 Fuzzy Sets
Below we provide a brief overview of fuzzy sets drawn from the work of Kruse et al. (2022). A fuzzy set $A$ is defined by its membership function $\mu_{A}:X\to[0,1]$, where $\mu_{A}(x)$ represents the degree of membership of $x\in X$ in $A$. The cardinality of $A$ is defined as the sum of the degrees of membership of all elements of $X$:
$\sum_{x\in X}\mu_{A}(x).$
If $A$ and $B$ are fuzzy sets over the same universe $X$ then their intersection is a fuzzy set whose membership function is defined as
$\mu_{A\cap B}(x)=t(\mu_{A}(x),\mu_{B}(x))$
[…]
A.2 Soft F1 Scores
Let $H_{k}$ and $R_{k}$ denote the predicted and the true (crisp) sets of examples in class $k$ respectively. Standard precision and recall scores for class $k$ are defined as:
$\mathrm{Prec}_{k}$ $=\frac{|H_{k}\cap R_{k}|}{|H_{k}|},$
$\mathrm{Rec}_{k}$ $=\frac{|H_{k}\cap R_{k}|}{|R_{k}|}.$
Embracing HLV, we view judgment distributions as degrees of memberships of fuzzy sets, each corresponds to a class, over a universe of examples. Specifically, we consider $P_{ik}$ and $Q_{ik}$ as the true and the predicted degrees of membership of example $i$ in class $k$. Therefore, we can define the soft precision and recall scores using fuzzy set operations as follows (Harju and Mesaros, 2023):
$\mathrm{SoftPrec}_{k}=\frac{\sum_{i}\min(P_{ik},Q_{ik})}{\sum_{i}Q_{ik}}$ (1)
and
$\mathrm{SoftRec}_{k}=\frac{\sum_{i}\min(P_{ik},Q_{ik})}{\sum_{i}P_{ik}}.$ (2)
The soft F1 score is the harmonic mean between the soft precision and recall scores as usual:
[…]
Taking the mean of Equation 3 over classes results in the soft macro F1 score. To obtain the micro variant, we simply modify the sums in both the numerator and the denominator of Equations 1 and 2 to also iterate over classes to get the soft micro precision and recall scores, and then take the harmonic mean of the two scores as normal.
Appendix B Experimental Setup of Pairwise Correlation
For a given number of examples and classes, we sample $B$ matrices $\mathbf{P}$ and $\mathbf{Q}$ and compute the value of each metric on these $B$ pairs of $\mathbf{P}$ and $\mathbf{Q}$. The Pearson correlations between pairs of metrics are then computed based on these $B$ data points. We set $B=500$ for computational reasons. We report the correlation coefficients in Table 2.
Appendix C Proof of Upper-Boundedness of Soft Accuracy by PO-JSD
Theorem 1.
Consider the same definitions of $P_{ik}$ and $Q_{ik}$ used in Section 3. The soft accuracy and the PO-JSD between $\mathbf{P}$ and $\mathbf{Q}$ satisfy the following inequality:
$\frac{1}{N}\sum_{ik}\min(P_{ik},Q_{ik})\leq 1-\frac{1}{N}\sum_{i}\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i})$
where $\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i})$ denote the Jensen-Shannon divergence between the $i$-th rows of $\mathbf{P}$ and $\mathbf{Q}$.
Proof.
Given a scalar $0\leq\pi\leq 1$ and discrete distributions $\mathbf{a},\mathbf{b}$, it has been shown that the following bound holds (Lin, 1991, Theorem 4):
$\sum_{k}\min(\pi a_{k},(1-\pi)b_{k})\\ \leq\frac{1}{2}\left(H([\pi,1-\pi])-\operatorname{JSD}_{\pi}(\mathbf{a},\mathbf{b})\right)$
where $H(\mathbf{u})$ is the entropy of $\mathbf{u}$, and $\operatorname{JSD}_{\pi}(\mathbf{a},\mathbf{b})=H(\pi\mathbf{a}+(1-\pi)\mathbf{b})-\pi H(\mathbf{a})-(1-\pi)H(\mathbf{b})$ is the general form of Jensen-Shannon divergence with two distributions. It can be easily shown that $\operatorname{JSD}_{\pi}$ is reduced to the standard Jensen-Shannon divergence when $\pi=\frac{1}{2}$. Performing this substitution and noting that $H([\frac{1}{2},\frac{1}{2}])=1$, [^18] All logarithms are of base 2. we have
$\sum_{k}\min(\frac{1}{2}P_{ik},\frac{1}{2}Q_{ik})\leq\frac{1}{2}\left(1-\operatorname{JSD}(\mathbf{p}_{i},\mathbf{q}_{i})\right)$
for all $i$. Doubling both sides and taking the average of this inequality over $i$ give the desired result. ∎
Dataset Model Acc./Micro F1 Macro F1 Entropy corr. PO-JSD Soft acc./micro F1 Soft macro F1
HS-Brexit TwHIN-BERT .903 $\pm$ .009 .581 $\pm$ .093 .470 $\pm$ .119 .941 $\pm$ .010 .883 $\pm$ .013 .676 $\pm$ .071
LLaMA .950 $\pm$ .012 .660 $\pm$ .104 .330 $\pm$ .065 .939 $\pm$ .002 .886 $\pm$ .003 .623 $\pm$ .012
MD-Agreement TwHIN-BERT .809 $\pm$ .003 .779 $\pm$ .002 .379 $\pm$ .002 .921 $\pm$ .001 .811 $\pm$ .002 .790 $\pm$ .003
LLaMA .810 $\pm$ .001 .784 $\pm$ .001 .215 $\pm$ .019 .863 $\pm$ .003 .770 $\pm$ .002 .743 $\pm$ .003
ArMIS TwHIN-BERT .713 $\pm$ .014 .705 $\pm$ .014 .062 $\pm$ .015 .769 $\pm$ .011 .702 $\pm$ .009 .695 $\pm$ .008
C-SNLI RoBERTa .632 $\pm$ .005 .551 $\pm$ .007 .123 $\pm$ .014 .830 $\pm$ .005 .672 $\pm$ .004 .629 $\pm$ .007
LLaMA .601 $\pm$ .010 .551 $\pm$ .008 .118 $\pm$ .022 .830 $\pm$ .005 .668 $\pm$ .005 .633 $\pm$ .007
C-MNLI RoBERTa .496 $\pm$ .014 .369 $\pm$ .014 .044 $\pm$ .007 .854 $\pm$ .009 .668 $\pm$ .010 .642 $\pm$ .009
LLaMA .514 $\pm$ .017 .439 $\pm$ .010 .100 $\pm$ .018 .827 $\pm$ .005 .643 $\pm$ .006 .615 $\pm$ .006
MFRC RoBERTa .657 $\pm$ .001 .456 $\pm$ .002 .404 $\pm$ .002 .743 $\pm$ .001 .591 $\pm$ .001 .418 $\pm$ .000
LLaMA .649 $\pm$ .001 .454 $\pm$ .005 .353 $\pm$ .005 .717 $\pm$ .001 .576 $\pm$ .001 .395 $\pm$ .001
TAG RoBERTa .742 $\pm$ .002 .610 $\pm$ .014 .394 $\pm$ .002 .985 $\pm$ .000 .645 $\pm$ .002 .471 $\pm$ .003
LLaMA .759 $\pm$ .006 .681 $\pm$ .005 .279 $\pm$ .010 .983 $\pm$ .000 .648 $\pm$ .002 .521 $\pm$ .003
Table 7: Mean ($\pm$ std) performance of MV measured by entropy correlation, PO-JSD, and various versions of F1 scores. C-SNLI and C-MNLI refer to the SNLI and the MNLI portions of ChaosNLI respectively. (Hard/soft) accuracy (for multiclass tasks) and (hard/soft) micro F1 score (for multilabel tasks) are reported in the same column.
(a) TwHIN-BERT on ArMIS
(b) The MNLI portion of ChaosNLI
Figure 4: Performance difference (delta) with mean MV performance on ArMIS (top) and the MNLI portion of ChaosNLI (bottom). LLaMA is omitted on ArMIS because its intended use is English while ArMIS is Arabic. C-MNLI doesn’t have annotator identity information, so AE, AEh, AR, ARh are inapplicable. The methods are sorted in the same manner as in Figure 2.
(a) Areas of law annotation
(b) Preference annotation
Figure 5: Annotation interface used to collect human preference data. Lawyers are asked to annotate areas of law for a given problem description (top) and then to select a judgment distribution is more accurate along with their certainty level (bottom).